OpenStudy (wyattp17):

what are these formulas in physics d=12gt sqaured F = ma d=m/v v=d/t F = G*m1*m2/d2

undeadknight26 (undeadknight26):

lol no idea...You would need one who expertise in this to know and remember these formulas.

OpenStudy (wyattp17):

tag one in this

OpenStudy (wyattp17):

@undeadknight26

OpenStudy (irishboy123):

d=12gt sqaured distance moved from rest in time t whilst undergoing constant acceleration a \[d = \frac{1}{2}a t^2\] F = ma force F = mass m times acceleration a, Newton's second d=m/v density = mass per unit volume v=d/t velocity = displacement (kinda like distance) divided by time F = G*m1*m2/d2 Newton's equation for gravitational attraction, F = force, G = universal gravitational constant, m1 m2 are masses of the bodies (eg planets), d is distance between them \[F = \frac{Gm_1m_2}{d^2}\]

OpenStudy (wyattp17):

thx irish @IrishBoy123